### v1

@@ Abstract
Depth sensing is an essential technology in robotics and many other fields. Many depth sensing (or RGB-D) cameras are available on the market and selecting the best one for your application can be challenging. In this work, we tested four stereoscopic RGB-D cameras that sense the distance by using two images from slightly different views. We empirically compared four cameras (Intel RealSense D435, Intel RealSense D455, StereoLabs ZED 2, and Luxonis OAK-D Pro) in three scenarios: (i) planar surface perception, (ii) plastic doll perception, (iii) household object perception (YCB dataset). We recorded and evaluated more than 3,000 RGB-D frames for each camera. For table-top robotics scenarios with distance to objects up to one meter, the best performance is provided by the D435 camera. For longer distances, the other three models perform better, making them more suitable for some mobile robotics applications. OAK-D Pro additionally offers integrated AI modules (e.g., object and human keypoint detection). ZED 2 is not a standalone device and requires a computer with a GPU for depth data acquisition. All data (more than 12,000 RGB-D frames) are publicly available at https://osf.io/f2seb.
@@ I Introduction
Depth sensing cameras, or RGB-D cameras, are visual sensors providing standard images (RGB), along with depth information. They are compact, powerful, and affordable sensors, making them attractive for a wide range of applications in robotics and many other fields. In this work, we compare four different RGB-D cameras: Intel RealSense D435 [1], Intel RealSense D455 [2], StereoLabs ZED 2 [3], and Luxonis OAK-D Pro [4]. All of the cameras work on the stereoscopic principle—the actual physical sensor contains two cameras with a fixed distance and the depth is computed from matching features in the two streams. RealSense and OAK-D Pro devices are active, i.e., they project an IR grid that helps with feature matching. Depth can be also perceived using Time-of-Flight (ToF) sensors, like LIDAR, which perceive depth more accurately, but are more bulky and expensive. They are typically used on large mobile robots and are not deployed in indoor robot manipulation [5].
@@
The accuracy of the sensors is crucial for the correct use and deployment. However, manufacturers usually provide only technical specifications and information on the percentual accuracy in a given depth [1, 2, 4, 3]. This information is useful, but it is not informative about the standard deviation in individual underlying tasks and was usually measured under unknown laboratory conditions that are hard to replicate in daily use. Researchers always wanted to better estimate the camera capabilities in computer vision applications, as done by et al.Andersen [6] with the Kinect or et al.Nguyen [7] who tried to estimate the noise of the Kinect to improve its performance in Simultaneous Localization and Mapping (SLAM). After Kinect, Intel cameras became popular. et al.Carfagni provided a performance evaluation of Intel RS300 [8]. Some researchers also compared stereo vison cameras with ToF cameras [9] or only ToF [10]. et al.Tadic performed comparison of Intel and StereoLabs cameras from the point-of-view of technical specifications and from visual inspection of depth images [11]. Other ZED camera analysis was performed in [12], where the authors tested the camera in indoor environments. Halmetschlager et al. [13] compared 10 depth-sensing cameras with different working principles, proposed a noise model, and advised which camera to use for which purpose.
@@
Some works focused more on accuracy in a given task, e.g., bolt tightening [14] or fruit detection and sizing [15, 16]. More unusual research containing RGB-D cameras can be the comparison with multi-view stereo for face surface reconstruction [17] or learning how to recognize a camera model or light source only from depth noise [18]. Similar comparisons to ours were made by et al.Heinemann [19] who compared six cameras, including the ones we have, in bias and standard deviation on flat surfaces, spheres and one 3D box. Other recent and comparable study was proposed by et al.Servi [20]. The authors compared 4 cameras (not the same set as us) using planes, spheres, and a 3D printed statue.
@@ II Cameras and Specifications
This section describes the hardware used and its technical specifications. The RGB-D cameras used in this work can be seen in Fig. 1 and their main technical specifications can be found in Table I. All four cameras use the stereoscopic effect to compute the depth. In other words, the depth is computed from correspondences in two images from slightly different views. Both RealSense devices and the OAK-D Pro camera use active stereo, i.e., they project additional IR points that help to find correspondences.
@@
On-board computation. The Intel cameras (D435, D455) perform depth computation on-board, and resources of a host computer are used only for potential post-processing. ZED 2, on the other hand, requires the host computer to contain CUDA-enabled Graphical Processing Unit (GPU) to compute the depth. The benefit arising from GPU necessity is that StereoLabs provides a wide selection of Artifical Inteligence (AI) tools, such as object detection or body tracking, which can be run through their Application Programming Interface (API) directly using the camera output. The OAK-D Pro camera performs all computations on board, but still provides AI features that also run directly on the camera.
Depth accuracy. In terms of depth accuracy reported by the respective manufacturers, D455, ZED 2, and OAK-D Pro should perform similarly with an error of $<$ 2% at 4 m. The D435 should, according to the specifications, be twice less accurate with error of $<$2% at 2 m. The devices that do not require GPU (D435, D455, OAK-D Pro) have similar maximal depth resolution (1280x720). However, the Frames per Second (FPS) with the maximal resolution is 30 for D435, D455 and 120 for OAK-D Pro. RealSense devices can get up to 90, but the resolution needs to be lower. These three cameras also have a similar Field of View (FOV). For ZED 2, the maximum resolution can be up to 2K with 15FPS (ZED 2 can get up to 100FPS based on the resolution) with larger FOV.
@@ III Setup and Evaluation
The setup utilized to record the experiments is depicted in Fig. 2. We mounted the four cameras on tripods in a long hallway and set the tripods so that the sensors of the individual sensors were at the same height and level. In front of the cameras, there is a table. We always moved the table or objects on top of the table, not the cameras. We tested three scenarios: (i) planar surface perception, where only the planar surface behind the doll in Fig. 2 was recorded; (ii) plastic doll perception, where a doll was captured by the cameras (corresponds to Fig. 2); (iii) and YCB objects perception, where objects from the YCB Object and Model set  [21] https://www.ycbbenchmarks.com/ were perceived (without the planar surface behind). The scenarios were selected to compare the cameras in terms of flat surface sensing (planes) and on more complicated objects (doll, YCB) as it is an important task for underlying robotic tasks such as segmentation, perception, or grasping.
The cameras were connected to one computer with a NVIDIA RTX 3070 GPU and Intel Core i9-11900 CPU. To eliminate the computational processing requirements, we always recorded output from one camera at a time. RealSense cameras were recorded through the RealSense Viewer application. ZED 2 and OAK-D Pro through their respective Python API. We used default settings for each camera—see Table II. No post-processing was applied.
@@ III-A Data Acquisition and Processing
  • Plastic doll and YCB objects – In this case, we moved the table further away and also moved the individual objects so that they are in front of the respective camera all the time. Then, from the resulting point clouds, we segmented the object. The segmentation was partially automatic (plane extraction, depth cropping) and partially manual to properly segment only the object of interest.
To evaluate the results, we needed to fit the segmented point clouds to a ground truth. We utilized the Iterative Closest Point (ICP[22] algorithm for an initial transformation. Further, we manually checked every estimated transformation and refined it to fit the ground truth properly. We used more than one frame for each object (for each camera at every depth), but we used only one transformation for the given setting. Otherwise, ICP could change the accuracy results, as it could mitigate the error by a transformation.
@@ III-B Ground-truth Point Clouds
For some metrics, ground-truth point clouds (shapes) are needed. For planar surfaces, we simply created a point cloud of a square with a side of 10 cm. For the doll, there is no ground truth available. We utilized photogrammetry to create the model of the doll, i.e., we took a number of pictures of the doll from various angles and used Meshroom [23] to create a point cloud of the doll. The YCB dataset was selected because it provides real-world items with corresponding models of the objects. Thus, for objects from YCB we downloaded the corresponding model and sampled them into point clouds. For the transparent spray bottle, ground truth is not available from YCB, so we used photogrammetry like in the case of the plastic doll.
The ground truth point clouds for the YCB objects are complete point clouds from all sides. However, in our experiments, we obtained point clouds from one view only. To make comparison fair, we manually cropped the ground-truth point clouds to match the viewpoint of the depth cameras.
@@ IV Experiments and Results
We tested the four cameras in three different scenarios: (i) planar surface perception, (ii) plastic doll perception, (iii) YCB objects perception. We measured each of these at several distances. For every object experiment, we used 30 successive frames, converted them to point clouds, and segmented the object out of the frame. Distances from the camera to the objects were measured with a laser measuring device and data were captured under constant light conditions of about 420 lx. In total, we compared the cameras with more than 3,000 point clouds from each camera, resulting in more than 12,000 total point clouds. All segmented point clouds, together with depth and RGB images, can be found at https://osf.io/f2seb.
@@ IV-A Planar Surface Perception
For standard deviation of distance from 30 measurements, all cameras behave similarly up to 100 cm with error $<$0.5 cm. Then, the RealSense devices (D435, D455) start to lose precision; this is also the case for the OAK-D Pro but to a lesser extent. In case of ZED 2, the standard deviation is $<$1.5 cm for all distances.
@@
The standard deviation can also be seen in the right columns of Fig. 5, where it is represented as the scatter of the points on the sides. It is well visible that, except for ZED 2, the width of the planes is getting bigger with increasing distance, corresponding to a increase in standard deviation in Fig. 4.
Another metric usable for planes perception is the angle between the closest normals in camera and ground-truth point clouds. In case of planes, this metric is basically a metric of flatness. The results can be seen in Fig. 6. Up to 80 cm all cameras stay under 0.2 rad (~10). However, the RealSense devices have twice the error even at the lower distances. At higher distances, the error of the RealSense devices increases to around 1 rad (~60°60°60  °). ZED 2 stays $<$0.2 rad and OAK-D Pro $<$0.1 rad for all distances. Farther than 100 cm from the camera, The RealSense devices (D435, D455) have more variability over different frames, while OAK-D Pro and ZED 2 show only small variability. The planes extracted from D435 and D455 are more uneven with increasing distance. The planes extracted from ZED 2 and OAK-D Pro are equally flat for all distances. OAK-D Pro, in fact, always returns points in almost “flat layers”, and thus the error from a flat surface is low.
@@
Planar surface perception – summary. All the results, Fig. 4, Fig. 5, Fig. 6 paint a coherent picture. For distances above one meter, D435 and D455 start losing both accuracy (positive bias – distance overestimation) and precision (greater variability). ZED 2 has very good performance even for the maximum tested distance (400 cm).
@@ IV-B Plastic Doll Perception
The second set of experiments involved perception of a plastic doll shown in Fig. 2. We recorded the samples at 90, 100, 125, 150, 175, 200, 250, 300, 350, and 400 cm from the camera. Point clouds taken from 125 cm and 400 cm can be seen in Fig. 7. We can see that at the closer distance, the point clouds look visually fine from both front and side views. The most visible flaws are at the forehead, which is a challenging part for RGB-D cameras as it is featureless and shiny. On the other hand, in the farther distance the point clouds are much noisier. The noise is visible especially from the side view, where the uneven depth estimation is visible for all cameras. However, it is interesting how different cameras handle the depth differently. RealSense devices divide the depth into layers with small differences between them and include some noisy points. OAK-D Pro divides the distances into layers with higher distances without visible noise. ZED 2 has the lowest depth difference among individual frames, but has the most noisy points.
@@
The other metric is the angle between normals. Here, it is not a measure of flatness as in the case of planar surfaces, but rather a measure of shape estimation. The results can be seen in Fig. 9. An important thing to notice is the high error of the OAK-D Pro camera at all depths. As has been said, OAK-D Pro creates layers of points at different depths (the depth differences between layers are lower for a lower object distance from the camera) instead of a scatter distribution, as is for other cameras. That results in facing of the normals more in the direction of the camera plane. Among the other cameras, ZED 2 and D435 are on par for every distance and D455 is worse up to 150 cm (again, we attribute this to the wider baseline), and then all cameras perform almost the same. Overall, the angles are higher than for the planar surfaces in lower distances—around 0.2 rad vs. 0.4 rad at the closest distance. However, it is interesting to look at the higher distances. RealSense devices achieve the same error as for the planes—about 1 rad. Both ZED 2 and OAK-D Pro are on par with Intel devices, but in the case of planes the error for the highest distance was under 0.2 rad for ZED 2 and 0.1 rad for OAK-D Pro. That is a huge difference that shows how the devices sense simple (planar) and more complex environments differently.
@@
For the final two metrics— Jaccard Similarity (JS) and $F_{1}$ score—the higher the value, the better. Both of these metrics evaluate the similarity of the shapes. In case of JS, the output is “geometric closeness” of two shapes. For $F_{1}$, it is the probability that a point in a given point cloud should be where it is when considering the ground truth. For D435, D455, and OAK-D Pro, the results are similar, with D435 being the best among those. In case of JS, the similarity is getting lower with a steeper drop from 250 cm further. $F_{1}$ similarity also demonstrates the same drop. For ZED 2, JS is getting gradually lower without any significant drop, and $F_{1}$ reveals only a slight drop (around 0.1) at higher distances. The reason is the amount of noise. As can be seen in the bottom block of Fig. 7, the ZED 2 point clouds have the lowest variance in depth estimates between frames, and the noise is more concentrated around the true values, compared to other cameras that produce more scattered results.
Plastic doll perception – summary. All the cameras performs similarly well up to 150 cm in all metrics. The best overall cameras is the ZED 2, that achieves CD error $<3\;cm$ even at the farthest distance (400 cm). The worst camera in this task is the OAK-D Pro. Both Intel devices perform similarly and up to 200 cm they are comparable to ZED 2 camera.

### v2

@@ Abstract
Depth sensing is an essential technology in robotics and many other fields. Many depth sensing (or RGB-D) cameras are available on the market and selecting the best one for your application can be challenging. In this work, we tested four stereoscopic RGB-D cameras that sense the distance by using two images from slightly different views. We empirically compared four cameras (Intel RealSense D435, Intel RealSense D455, StereoLabs ZED 2, and Luxonis OAK-D Pro) in three scenarios: (i) planar surface perception, (ii) plastic doll perception, (iii) household object perception (YCB dataset). We recorded and evaluated more than 3,000 RGB-D frames for each camera. For table-top robotics scenarios with distance to objects up to one meter, the best performance is provided by the D435 camera that is able to perceive with an error under 1 cm in all of the tested scenarios. For longer distances, the other three models perform better, making them more suitable for some mobile robotics applications. OAK-D Pro additionally offers integrated AI modules (e.g., object and human keypoint detection). ZED 2 is overall the best camera which is able to keep the error under 3 cm even at 4 meters. However, it is not a standalone device and requires a computer with a GPU for depth data acquisition. All data (more than 12,000 RGB-D frames) are publicly available at https://rustlluk.github.io/rgbd-comparison.
@@ I Introduction
Depth sensing cameras, or RGB-D cameras, are visual sensors that provide standard images (RGB), along with depth information. They are compact, powerful, and affordable sensors which make them attractive for a wide range of applications in robotics and many other fields. Depth information is useful when building a map of an environment, so RGB-D cameras are widely used in Simultaneous Localization and Mapping (SLAM[1, 2] or scene reconstruction [3, 4, 5]. They can also be used for 3D reconstruction or shape completion of individual objects [6, 7]. An important table-top task that utilizes depth sensing is grasping [8, 9, 10]. An interesting application is human pose and shape estimation [11, 12, 13]. The output can then be used, for example, in human-robot interaction [14, 13].
In this work, we compare four different RGB-D cameras: Intel RealSense D435 [15], Intel RealSense D455 [16], StereoLabs ZED 2 [17], and Luxonis OAK-D Pro [18]. RGB-D cameras have many applications. However, currently the state of the art in camera comparison is usually focusing only on performance on flat surfaces, which is not representative of the different application scenarios. The goals of this work are thus two-fold. First, we want to complement the performance comparison on planar surfaces with the perception of 3D objects with complicated geometries, focusing also on shorter distances relevant for robot manipulation, for example. Second, the set of cameras we tested is new and aims to reflect the most relevant models available on the market now.
@@
All of the cameras selected in this work operate on the stereoscopic principle—the actual physical sensor contains two cameras with a fixed distance, and the depth is computed from matching features in the two streams. RealSense and OAK-D Pro devices are active, i.e., they project an IR grid that helps with feature matching. Depth can be also perceived using Time-of-Flight (ToF) sensors, like LIDAR, which perceive depth more accurately, but are more bulky and expensive. They are typically used on large mobile robots and are not deployed in indoor robot manipulation [19].
@@
The accuracy of the sensors is crucial for the correct use and deployment. However, manufacturers usually provide only technical specifications and information on the percentual accuracy in a given depth [15, 16, 18, 17]. This information is useful, but it is not informative about the standard deviation in individual underlying tasks and was usually measured under unknown laboratory conditions that are hard to replicate in daily use. Researchers always wanted to better estimate the camera capabilities in computer vision applications, as done by et al.Andersen [20] with the Kinect or et al.Nguyen [21] who tried to estimate the noise of the Kinect to improve its performance in SLAM. After Kinect, Intel cameras became popular. et al.Carfagni provided a performance evaluation of Intel RS300 [22]. Some researchers also compared stereo vison cameras with ToF cameras [23] or only ToF [24]. et al.Tadic performed comparison of Intel and StereoLabs cameras from the point-of-view of technical specifications and from visual inspection of the depth images [25]. Other ZED camera analysis was performed in [26], where the authors tested the camera in indoor environments. Halmetschlager et al. [27] compared 10 depth-sensing cameras with different working principles, proposed a noise model, and advised which camera to use for which purpose.
@@
Some works focused more on accuracy in a given task, e.g., bolt tightening [28] or fruit detection and sizing [29, 30]. More unusual research containing RGB-D cameras can be the comparison with multi-view stereo for face surface reconstruction [31] or learning how to recognize a camera model or light source only from depth noise [32]. Similar comparisons to ours were made by et al.Heinemann [33] who compared six cameras, including the ones we have, in bias and standard deviation on flat surfaces, spheres, and one 3D box. Other recent and comparable study was proposed by et al.Servi [34]. The authors compared 4 cameras (not the same set as us) using planes, spheres, and a 3D printed statue.
@@ II Cameras and Specifications
This section describes the hardware used and its technical specifications. The RGB-D cameras used in this work can be seen in Fig. 1 and their main technical specifications can be found in Table I. The selection was motivated by an informal survey through the recent literature and among colleagues in robotics (mobile robotics and robot manipulation). All four cameras use the stereoscopic effect to compute the depth. In other words, the depth is computed from correspondences in two images from slightly different views. Both RealSense devices and the OAK-D Pro camera use active stereo, i.e., they project additional IR points that help to find correspondences.
@@
On-board computation. The Intel cameras (D435, D455) perform depth computation on-board, and the resources of a host computer are used only for potential post-processing. ZED 2, on the other hand, requires the host computer to contain CUDA-enabled Graphical Processing Unit (GPU) to compute the depth. The benefit arising from GPU necessity is that StereoLabs provides a wide selection of Artifical Inteligence (AI) tools, such as object detection or body tracking, which can be run through their Application Programming Interface (API) directly using camera output. The OAK-D Pro camera performs all computations on board, but still provides AI features that also run directly on the camera.
Depth accuracy. In terms of depth accuracy reported by the respective manufacturers, D455, ZED 2, and OAK-D Pro should perform similarly with an error of $<$ 2% at 4 m. The D435 should, according to the specifications, be twice less accurate with error of $<$2% at 2 m. The devices that do not require GPU (D435, D455, OAK-D Pro) have a similar maximal depth resolution (1280x720). However, the Frames per Second (FPS) with the maximal resolution is 30 for D435, D455 and 120 for OAK-D Pro. RealSense devices can get up to 90, but the resolution needs to be lower. These three cameras also have a similar Field of View (FOV). For ZED 2, the maximum resolution can be up to 2K with 15FPS (ZED 2 can get up to 100FPS based on the resolution) with larger FOV.
@@ III Setup and Evaluation
The setup utilized to record the experiments is depicted in Fig. 2. We mounted the four cameras on tripods in a long hallway and set the tripods so that the sensors of the individual sensors were at the same height and level. The distance was measured using a laser measuring device from a plane surface. The same height was achieved using a long spirit level. In front of the cameras, there is a table. We always moved the table or objects on top of the table, not the cameras. We tested three scenarios: (i) planar surface perception, where only the planar surface behind the doll in Fig. 2 was recorded; (ii) plastic doll perception, where a doll was captured by the cameras (corresponds to Fig. 2); (iii) and YCB objects perception, where objects from the YCB Object and Model set  [35] https://www.ycbbenchmarks.com/ were perceived (without the planar surface behind).
The scenarios were selected to compare the cameras in terms of flat surface sensing (planes) and on more complicated objects (doll, YCB). Flat surfaces are important in mobile robotics (e.g., navigation in indoor environments) or to easily segment tables in table-top scenarios. The plastic doll was selected as a complicated 3D object and at the same time a human figure, as RGB-D cameras are also used in human pose estimation and shape reconstruction scenarios. The YCB dataset is widely used for benchmarking in object segmentation and robot manipulation; the video variant YCB-Video [36] is used in 6D pose estimation.
The distances from the camera to the objects were measured with a laser measuring device and the data were captured under constant light conditions of about 420 lx. As we recorded in a long hallway, we did not have a proper way to control the light conditions. Thus, to keep the comparison as fair and as consistent as possible, we switched on the light (which were equally distributed over the distance) and checked the light intensity before recording at every distance using a lux meter.
The cameras were connected to one computer with a NVIDIA RTX 3070 GPU and Intel Core i9-11900 CPU. To eliminate computational processing requirements, we always recorded the output from one camera at a time. RealSense cameras were recorded through the RealSense Viewer application. ZED 2 and OAK-D Pro through their respective Python API. We used default settings for each camera—see Table II. No post-processing was applied.
@@ III-A Data Acquisition and Processing
  • Plastic doll and YCB objects – In this case, we moved the table further away and also moved the individual objects so that they are in front of the respective camera all the time. Then, from the resulting point clouds, we segmented the object. We cropped the point clouds in such a way to keep only a region around the distance at which we recorded the object. At lower distances, we further utilized the RANSAC [37] algorithm to automatically extract the largest plane—the table—and cropped everything “under” it to get the segmented object. At farther distances, this procedure did not work as the plane was not flat anymore and thus RANSAC was unable to find the correct plane. Therefore, we manually segmented the object from the surroundings. Ultimately, we checked all individual point clouds for correct segmentation.
To evaluate the results, we needed to fit the segmented point clouds to a ground truth. We utilized the Iterative Closest Point (ICP[38] algorithm for an initial transformation. The accuracy of ICP is limited when used only on partial point clouds of relatively small size (as in our case). Thus, for each captured and ground-truth point cloud we ran the algorithm with different convergence and inliers parameters and selected the best one as the one with the lowest distance between the estimated corresponding points. However, as this may still do not guarantee correct results at all times—especially at farther distances, where the objects are small and noisy. Thus, we further manually checked every estimated transformation and refined it to fit the ground truth properly. We used more than one frame for each object (for each camera at every depth), but we used only one transformation for the given setting. Otherwise, ICP could change the accuracy results, as it could mitigate the error by a transformation.
@@ III-B Ground-truth Point Clouds
For some metrics, ground-truth point clouds (shapes) are needed. For planar surfaces, we simply created a point cloud of a square with a side of 10 cm. For the doll, there is no ground truth available. We utilized photogrammetry to create the model of the doll using Meshroom [39]. The process consisted of taking pictures of the doll from various angles and view-points. Meshroom then performed feature matching using Scale-Invariant Feature Transform (SIFT) from the unordered set of images. The features were then used to match pictures looking at the same areas of the scene and Structure from Motion (SfM) algorithm was used to create a 2D point cloud of the scene. Further, fronto-parallel planes were used to estimate depth for each of the pixels from SfM. Finally, the resulting 3D point cloud was meshed using Delaunay tetrahedralization and textured using the input image. To add scale to the model, we put markers beside the doll with a known distance and we then scaled the model accordingly. We then manually refined the final 3D triangular mesh to remove unwanted holes and artifacts.
Figure 3: Objects from the YCB dataset used in the experiments.
The YCB dataset was selected because it provides real-world items with corresponding models of the objects—see the used objects in Fig. 3. Thus, for objects from YCB we downloaded the corresponding model and sampled them into point clouds. For the transparent spray bottle, ground truth is not available from YCB, so we used photogrammetry as in the case of the plastic doll.
The ground truth point clouds for the YCB objects are complete point clouds from all sides. However, in our experiments, we obtained point clouds from one view only. To make the comparison fair, we manually cropped the ground-truth point clouds to match the viewpoint of the depth cameras.
@@ IV Experiments and Results
We tested the four cameras in three different scenarios: (i) planar surface perception, (ii) plastic doll perception, (iii) YCB objects perception. We measured each of these at several distances. For every object experiment, we used 30 successive frames, converted them to point clouds, and segmented the object out of the frame. In total, we compared the cameras with more than 3,000 point clouds from each camera, resulting in more than 12,000 total point clouds. All segmented point clouds, together with depth and RGB images, can be found at https://rustlluk.github.io/rgbd-comparison.
@@ IV-A Planar Surface Perception
For the standard deviation of distance from 30 measurements, all cameras behave similarly up to 100 cm with error $<$0.5 cm. Then, the RealSense devices (D435, D455) start to lose precision; this is also the case for the OAK-D Pro but to a lesser extent. In case of ZED 2, the standard deviation is $<$1.5 cm for all distances. We can also compare with the accuracy provided by the manufacturers. The comparison is shown in Table III. It is not always known what the term “accuracy” means for each of the cameras, but we assume that it will be comparable to either bias or standard deviation. We can see that the real errors are lower than the ones reported by the manufacturers, except for the standard deviation of D455.
TABLE III: Planar surface perception – comparison with factory accuracy.
Factory accuracy Standard deviation Bias
D435 $<$2% at 2 m $0.75$% at 2 m $1.87$% at 2 m
D455 $<$2% at 4 m $3.05$% at 4 m $0.13$% at 4 m
ZED $<$0.8% at 2 m $0.17$% at 2 m $0.22$% at 2 m
OAK-D Pro $<$2% at 4 m $1.47$% at 4 m $1.65$% at 4 m
@@
The standard deviation can also be seen in the right columns of Fig. 6, where it is represented as the scatter of the points on the sides. It is well visible that, except for ZED 2, the width of the planes increases with increasing distance, corresponding to an increase in standard deviation in Fig. 5.
@@
Another metric usable for plane perception is the angle between the closest normals in camera and ground-truth point clouds. In case of planes, this metric is basically a metric of flatness. The results can be seen in Fig. 7. Up to 80 cm all cameras stay under 0.2 rad (~10). However, the RealSense devices have twice the error even at the lower distances. At higher distances, the error of the RealSense devices increases to around 1 rad (~60°60°60  °). ZED 2 stays $<$0.2 rad and OAK-D Pro $<$0.1 rad for all distances. Farther than 100 cm from the camera, the RealSense devices (D435, D455) have more variability over different frames, while OAK-D Pro and ZED 2 show only small variability. The planes extracted from D435 and D455 are more uneven with increasing distance. The planes extracted from ZED 2 and OAK-D Pro are equally flat for all distances. In fact, OAK-D Pro always returns points in almost “flat layers”, and thus the error from a flat surface is low.
Planar surface perception – summary. All the results Fig. 5, Fig. 6, Fig. 7 paint a coherent picture. For distances above one meter, D435 and D455 start losing both accuracy (positive bias – distance overestimation) and precision (greater variability). ZED 2 has very good performance even for the maximum tested distance (400 cm). Planar surface perception is an important task for mobile navigation, mainly in indoor environments. Based on the results, the best choice in this scenario would be D435, OAK-D Pro, or ZED 2. The choice between these is then mainly related to the fact whether you can put a dedicated GPU on the robot or not.
@@ IV-B Plastic Doll Perception
The second set of experiments involved the perception of a plastic doll shown in Fig. 2. We recorded the samples at 90, 100, 125, 150, 175, 200, 250, 300, 350, and 400 cm from the camera. Point clouds taken from 125 cm and 400 cm can be seen in Fig. 8. We can see that at the closer distance, the point clouds look visually fine from both front and side views. The most visible flaws are at the forehead, which is a challenging part for RGB-D cameras as it is featureless and shiny. On the other hand, in the farther distance the point clouds are much noisier. The noise is visible especially from the side view, where the uneven depth estimation is visible for all cameras. However, it is interesting how different cameras handle the depth differently. RealSense devices divide the depth into layers with small differences between them and include some noisy points. OAK-D Pro divides the distances into layers with higher distances without visible noise. ZED 2 has the lowest depth difference among individual frames, but has the most noisy points.
@@
The other metric is the angle between normals. Here, it is not a measure of flatness as in the case of planar surfaces, but rather a measure of shape estimation. The results can be seen in Fig. 10. An important thing to notice is the high error of the OAK-D Pro camera at all depths. As has been said, OAK-D Pro creates layers of points at different depths (the depth differences between layers are lower for a lower object distance from the camera) instead of a scatter distribution, as is for other cameras. That results in facing of the normals more in the direction of the camera plane. Among the other cameras, ZED 2 and D435 are on par for every distance, and D455 is worse up to 150 cm (again, we attribute this to the wider baseline), and then all cameras perform almost the same. Overall, the angles are higher than for the planar surfaces in lower distances—around 0.2 rad vs. 0.4 rad at the closest distance. However, it is interesting to look at the higher distances. RealSense devices achieve the same error as for the planes—about 1 rad. Both ZED 2 and OAK-D Pro are on par with Intel devices, but in the case of planes the error for the highest distance was under 0.2 rad for ZED 2 and 0.1 rad for OAK-D Pro. That is a huge difference that shows how the devices sense simple (planar) and more complex environments differently.
@@
For the final two metrics— Jaccard Similarity (JS) and $F_{1}$ score—the higher the value, the better. Both of these metrics evaluate the similarity of the shapes. In case of JS, the output is “geometric closeness” of two shapes. For $F_{1}$, it is the probability that a point in a given point cloud should be where it is when considering the ground truth. For D435, D455, and OAK-D Pro, the results are similar, with D435 being the best among these. In case of JS, the similarity is decreasing with a steeper drop from 250 cm further. $F_{1}$ similarity also demonstrates the same drop. For ZED 2, JS is getting gradually lower without any significant drop, and $F_{1}$ reveals only a slight drop (around 0.1) at higher distances. The reason is the amount of noise. As can be seen in the bottom block of Fig. 8, the ZED 2 point clouds have the lowest variance in depth estimates between frames, and the noise is more concentrated around the true values, compared to other cameras that produce more scattered results.
Plastic doll perception – summary. All the cameras performs similarly well up to 150 cm in all metrics. The best overall cameras is the ZED 2, that achieves CD error $<3\;cm$ even at the farthest distance (400 cm). The worst camera in this task is the OAK-D Pro. Both Intel devices perform similarly and up to 200 cm they are comparable to the ZED 2 camera. This scenario simulates mainly the task of human pose and shape estimation. If humans are perceived from a close distance, typical for human-robot interaction scenarios, it is better to choose either D435 or ZED 2. ZED 2 directly offers keypoint detection algorithms.